OpenStudy (anonymous):

what is the derivative of 4/x what is the derivative of 4/x @Calculus1

OpenStudy (anonymous):

\[\frac{d}{dx}\frac{4}{x}=4*\frac{d}{dx}\frac{1}{x}= -\frac{4}{x^2}\]

OpenStudy (anonymous):

well the full question is on the other question. it's, locate the absolute extrema of the function on the closed interval